### v1

# Rhaly operators acting on Hardy and Bergman spaces

Abstract.
In this article we address the question of characterizing the sequences of complex numbers $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ whose associated Rhaly operator $\mathcal{R}_{(\eta)}$ is bounded or compact on the Hardy spaces $H^{p}$ ($1\leq p<\infty$) or on the Bergman spaces $A^{p}$ ($1\leq p<\infty$). Among other results we completely characterize those $(\eta)$ for which $\mathcal{R}_{(\eta)}$ is bounded or compact on $H^{p}$ ($1<p\leq 2$) and on $A^{p}$ ($1<p<\infty$). We also give conditions on $(\eta)$ which are either necessary or sufficient for the boundedness (compactness) of $\mathcal{R}_{(\eta)}$ on $H^{p}$ for $p=1$ and $2<p<\infty$.
In particular, we prove that if $2\leq p<\infty$ and $\eta_{n}=\operatorname{O}\left(\frac{1}{n}\right)$, then $\mathcal{R}_{(\eta)}$ is bounded on $H^{p}$. However, there exists a sequence $(\eta)$ with $\eta_{n}=\operatorname{O}\left(\frac{1}{n}\right)$ such that the operator $\mathcal{R}_{(\eta)}$ is not bounded on $H^{p}$ for $1\leq p<2$.
Key words and phrases:
Hardy spaces, Bergman spaces, The Cesàro operator, Rhaly operators, Mean Lipschitz spaces
2020 Mathematics Subject Classification:
Primary 30H10; 47B91; 42B30
This research has been supported in part by a grant from “El Ministerio de Ciencia e Innovación”, Spain Project PID2022-133619NB-I00 and by grant from la Junta de Andalucía FQM-210-G-FEDER)
Corresponding author: Daniel Girela, girela@uma.es
1. Introduction and main results
If $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ is a sequence of complex numbers, the Rhaly matrix $\mathcal{R}_{(\eta)}=\mathcal{R}_{\{\eta_{n}\}}$ is defined by
$\mathcal{R}_{(\eta)}=\mathcal{R}_{\{\eta_{n}\}}\,=\,\left(\begin{array}[]{ccccccc}\eta_{0}&0&0&0&.&.\\ \eta_{1}&\eta_{1}&0&0&.&.\\ \eta_{2}&\eta_{2}&\eta_{2}&0&.&.\\ \eta_{3}&\eta_{3}&\eta_{3}&\eta_{3}&.&.\\ .&.&.&.&.&.\\ .&.&.&.&.&.\\ \end{array}\right).$
This matrix induces the Rhaly operator $\mathcal{R}_{(\eta)}=\mathcal{R}_{\{\eta_{n}\}}$ by matrix multiplication, that is, if $(a)=\{a_{n}\}_{n=0}^{\infty}$ is a sequence of complex numbers
$\mathcal{R}_{(\eta)}(a)\,=\,\mathcal{R}_{\{\eta_{n}\}}(a)\,=\,\left\{\eta_{n}\sum_{k=0}^{n}a_{k}\right\}_{n=0}^{\infty}.$
The case when $\eta_{n}=\frac{1}{n+1}$ for all $n$ corresponds to the classical Cesàro operator $\mathcal{C}$ which is known to be bounded from $\ell^{p}$ to $\ell^{p}$ for $1<p\leq\infty$. This was proved by Hardy [21] and Landau [25] (see also [22, Theorem 326, p.239 ]). The sequences $\{\eta_{n}\}_{n=0}^{\infty}$ for which the operator $\mathcal{R}_{\{\eta_{n}\}}$ is either bounded or compact on the space $\ell^{p}$ for any $p\in(1,\infty)$ have been characterized in [19]. A characterization of the sequences $\{\eta_{n}\}$ for which $\mathcal{R}_{\{\eta_{n}\}}$ belongs to the Schatten class $\mathcal{S}^{q}(\ell^{2})$ ($1<q<\infty$) has been obtained in [4].
Let ${\mathbb{D}}=\{z\in{\mathbb{C}}:|z|<1\}$ denote the open unit disc in the complex plane $\mathbb{C}$}. If $\mu$ is a complex Borel measure in $\mathbb{D}$ and, for $n=0,1,2,\dots$, $\mu_{n}$ denotes the moment of order $n$ of $\mu$, that is,
$\mu_{n}\,=\,\int_{\mathbb{D}}w^{n}\,d\mu(w),\quad n=0,1,2,\dots,$
the operator $\mathcal{R}_{\{\mu_{n}\}}$ will be denoted by $\mathcal{C}_{\mu}$. When $\mu$ is the Lebesgue measure on the radius $[0,1)$ the operator $\mathcal{C}_{\mu}$ is the Cesàro operator $\mathcal{C}$.
Let $\rm{Hol}({\mathbb{D}})$ be the space of all analytic functions in $\mathbb{D}$} endowed with the topology of uniform convergence in compact subsets.
If  0<r<1\,0<r<1\, and $f\in\rm{Hol}({\mathbb{D}})$, we set
$M_{p}(r,f)=\left(\frac{1}{2\pi}\int_{0}^{2\pi}|f(re^{it})|^{p}\,dt\right)^{1/p},\,\,\,0<p<\infty,$
$M_{\infty}(r,f)=\sup_{|z|=r}|f(z)|.$
For  $0<p\leq\infty$,  the Hardy space $H^{p}$ consists of those $f\in\rm{Hol}(\mathbb{D})$ such that
$\|f\|_{H^{p}}\stackrel{{\scriptstyle\text{def}}}{{=}}\sup_{0<r<1}M_{p}(r,f)<\infty.$
We refer to [11] for the notation and results regarding Hardy spaces.
Let $dA$ denote the area measure on $\mathbb{D}$, normalized so that the area of $\mathbb{D}$ is $1$. Thus $dA(z)\,=\,\frac{1}{\pi}\,dx\,dy\,=\,\frac{1}{\pi}\,r\,dr\,d\theta$. For $0<p<\infty$ and $\alpha>-1$ the weighted Bergman space $A^{p}_{\alpha}$ consists of those $f\in\rm{Hol}(\mathbb{D})$ such that
$\|f\|_{A^{p}_{\alpha}}\,\stackrel{{\scriptstyle\text{def}}}{{=}}\,\left(\int_{\mathbb{D}}|f(z)|^{p}\,dA_{\alpha}(z)\right)^{1/p}\,<\,\infty$,
where $dA_{\alpha}(z)=(\alpha+1)(1-|z|^{2})^{\alpha}dA(z)$. We refer to [12, 23, 43] for the notation and results about Bergman spaces.
If $0<p<\infty$ and $\alpha>-1$ the space of Dirichlet type $\mathcal{D}^{p}_{\alpha}$} consists of all primitives of functions in $A^{p}_{\alpha}$. Hence, if $f$ is analytic in $\mathbb{D}$, then $f\in{\mathcal{D}^{p}_{\alpha}}$ if and only if
$\|f\|_{{\mathcal{D}^{p}_{\alpha}}}^{p}\stackrel{{\scriptstyle\text{def}}}{{=}}|f(0)|^{p}+\|f^{\prime}\|_{A^{p}_{\alpha}}^{p}<\infty.$
Identifying any given function $f\in\rm{Hol}(\mathbb{D})$ with the sequence $\{a_{k}\}_{k=0}^{\infty}$ of its Taylor coefficients, the Cesàro operator $\mathcal{C}$ and the operators $\mathcal{C}_{\mu}$ associated to complex Borel measures on $\mathbb{D}$ become linear operators from $\rm{Hol}(\mathbb{D})$ into itself as follows:
If $f\in\rm{Hol}(\mathbb{D})$, $f(z)=\sum_{k=0}^{\infty}a_{k}z^{k}$  ($z\in\mathbb{D}$), then
$\mathcal{C}(f)(z)=\sum_{n=0}^{\infty}\left(\frac{1}{n+1}\sum_{k=0}^{n}a_{k}\right)z^{n},\quad\mathcal{C}_{\mu}(f)(z)=\sum_{n=0}^{\infty}\left(\mu_{n}\sum_{k=0}^{n}a_{k}\right)z^{n},\quad z\in\mathbb{D}.$
The Cesàro operator is known to be bounded on the Hardy spaces $H^{p}$ ($0<p<\infty$) and on the Bergman spaces $A^{p}_{\alpha}$ ($0<p<\infty$, $\alpha>-1$). This has been established by a number of authors using different methods [1, 31, 32, 35, 36, 37, 38].
The operators $\mathcal{C}_{\mu}$ with $\mu$ being a finite positive Borel measure supported on the radius $[0,1)$ were introduced in [15] where it was proved that such a $\mathcal{C}_{\mu}$ is bounded on the Hardy space $H^{p}$ ($1\leq p<\infty$), or on the Bergman space $A^{p}_{\alpha}$ ($1<p<\infty$, $\alpha>-1$) if and only if $\mu$ is a Carleson measure, that is, if there exists a positive constant $C$ such that $\mu([r,1))\,\leq\,C(1-r)$ for $0<r<1$ or, equivalently, if $\mu_{n}=\operatorname{O}\left(\frac{1}{n+1}\right)$. Theorem 9 of [14] shows that this is also true for the spaces $A^{1}_{\alpha}$ ($\alpha>-1$). Since its introduction, the boundedness of these operators $\mathcal{C}_{\mu}$ on other spaces of analytic functions in $\mathbb{D}$ has been studied by distinct authors (see, e. g., [3, 5, 17, 24, 39]).
Blasco [7] considered the operators $\mathcal{C}_{\mu}$, with $\mu$ being a complex Borel measure on $[0,1)$, acting on Hardy spaces. Later on, a good amount of work has been carried out considering the operators $\mathcal{C}_{\mu}$, where $\mu$ is a complex Borel measure on $\mathbb{D}$, acting on distinct subspaces of $\rm{Hol}(\mathbb{D})$; see, for instance, [6, 16, 26].
Let us remark that when $\mu$ is a complex Borel measure on $\mathbb{D}$, the operator $\mathcal{C}_{\mu}$ has an integral representation: If $f\in\rm{Hol}(\mathbb{D})$, $f(z)=\sum_{k=0}^{\infty}a_{k}z^{k}$ ($z\in\mathbb{D}$), then
(1.1) $\mathcal{C}_{\mu}(f)(z)\,=\,\sum_{n=0}^{\infty}\mu_{n}\left(\sum_{k=0}^{n}a_{k}\right)z^{n}\,=\,\int_{\mathbb{D}}\frac{f(wz)}{1-wz}\,d\mu(w),\quad z\in\mathbb{D}.$
This plays a crucial role to obtain a good number of the results we have just mentioned.
If $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ is a sequence of complex numbers then the Rhaly operator $\mathcal{R}_{\{\eta_{n}\}}$ can be viewed formally as an operator acting on spaces of analytic functions in the disc: If $f\in\rm{Hol}(\mathbb{D})$, $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ ($z\in\mathbb{D}$), formally, we set
$\mathcal{R}_{(\eta)}(f)(z)\,=\,\mathcal{R}_{\{\eta_{n}\}}(f)(z)=\sum_{n=0}^{\infty}\eta_{n}\left(\sum_{k=0}^{n}a_{k}\right)z^{n},\quad z\in\mathbb{D},$
whenever the right hand side makes sense and defines an analytic function in $\mathbb{D}$.
The sequence $(\eta)=\{\eta_{n}\}$ may not be bounded. Hence, in contrary to what happens with the operators $\mathcal{C}_{\mu}$, we cannot assure that the operator $\mathcal{R}_{\{\eta_{n}\}}$ is a well defined operator from $\rm{Hol}(\mathbb{D})$ into itself. Proposition 1 of [8] asserts that if $X$ is a subspace of $\rm{Hol}(\mathbb{D})$ which contains the constants, then $\mathcal{R}_{\{\eta_{n}\}}$ is a well defined linear operator from $X$ into $\rm{Hol}(\mathbb{D})$ if and only if the power series $\sum_{n=0}^{\infty}\eta_{n}z^{n}$ defines an analytic function in $\mathbb{D}$.
From now on, if the power series $\sum_{n=0}^{\infty}\eta_{n}z^{n}$ defines an analytic function in $\mathbb{D}$, we shall set
$F_{(\eta)}(z)\,=\,F_{\{\eta_{n}\}}(z)\,=\,\sum_{n=0}^{\infty}\eta_{n}z^{n},\quad z\in\mathbb{D}.$
In view of the identification of $H^{2}$ and $\ell^{2}$, the sequences $(\eta)$ for which $\mathcal{R}_{(\eta)}$ is bounded or compact on $H^{2}$ are characterized in [19]. The articles [2, 8, 14] contain characterizations of the sequences $(\eta)$ for which the operator $\mathcal{R}_{(\eta)}$ is bounded (compact) from $\mathcal{D}^{2}_{\alpha}$ into $\mathcal{D}^{2}_{\beta}$ ($\alpha,\beta\in\mathbb{R}$). Here, for $\alpha\in\mathbb{R}$, $\mathcal{D}^{2}_{\alpha}$ is the space of functions $f\in\rm{Hol}(\mathbb{D})$ such that $|a_{0}|^{2}+\sum_{n=1}^{\infty}n^{1-\alpha}|a_{n}|^{2}<\infty$, where $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$.
When we deal with a general Rhaly operator $\mathcal{R}_{(\eta)}$, we do not have an integral representation similar to (1.1). This is one reason why a characterization of those sequences $(\eta)$ for which $\mathcal{R}_{(\eta)}$ is bounded (compact) on $H^{p}$ or on $A^{p}_{\alpha}$ with $p\neq 2$ has not been obtained yet. In this paper we address the question of characterizing such sequences for $1<p<\infty$. For simplicity, if $X$ and $Y$ are two Banach spaces, $\mathcal{B}(X,Y)$ will denote the space of continuous linear operators from $X$ into $Y$ and $\mathcal{K}(X,Y)$ will stand for the space of all compact linear operators from $X$ into $Y$. Also $\mathcal{B}(X)$ will stand for $\mathcal{B}(X,X)$ and $\mathcal{K}(X)$ will stand for $\mathcal{K}(X,X)$. Our main results are the following.
Theorem 1.
Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers.
(a) If $1<p\leq 2$, then $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$  if and only if $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
(b) If $2<p<\infty$ and $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$, then $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
(c) If $2<p<\infty$ and $F_{(\eta)}\in\Lambda^{q}_{1/q}$ for some $q\in(2,p)$, then $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$.
Theorem 2.
Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers.
(a) If $1<p\leq 2$ then $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$  if and only if $F_{(\eta)}\in\lambda^{p}_{1/p}$.
(b) If $2<p<\infty$ and $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$, then $F_{(\eta)}\in\lambda^{p}_{1/p}$.
(c) If $2<p<\infty$ and $F_{(\eta)}\in\lambda^{q}_{1/q}$ for some $q\in(2,p)$, then $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$.
Theorem 3.
Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers and $1<p<\infty$. Then:
(a) $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p})$  if and only if $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
(b) $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p})$  if and only if $F_{(\eta)}\in\lambda^{p}_{1/p}$.
Since the Cesàro operator is bounded on $H^{p}$ for  $0<p<\infty$, it is natural to ask whether or not the condition $\eta_{n}\,=\,\operatorname{O}\left(\frac{1}{n}\right)$ implies that $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$ for some (all) $p$. We shall prove the following result.
Theorem 4.
(i) Suppose that $2\leq p<\infty$ and let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers. If  $\eta_{n}\,=\,\operatorname{O}\left(\frac{1}{n}\right)$, as $n\to\infty$, then $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$.
(ii) There exists a sequence of complex numbers $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$, with $\eta_{n}\,=\,\operatorname{O}\left(\frac{1}{n}\right)$, as $n\to\infty$, such that $\mathcal{R}_{(\eta)}\not\in\mathcal{B}(H^{p})$ for $1\leq p<2$.
(iii) If  $1<p<\infty$  and $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ is a decreasing sequence of non negative real numbers then $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$ if and only if $\eta_{n}\,=\,\operatorname{O}\left(\frac{1}{n}\right)$, as $n\to\infty$.
The mean Lipschitz spaces $\Lambda^{p}_{1/p}$ and the “litle oh”mean Lipschitz spaces $\lambda^{p}_{1/p}$ will be properly defined in Section 2, which will be devoted to introduce a number of spaces and results which will be needed in our work. The proofs of Theorem 1 and Theorem 2 will be presented in Section 3, and Section 4 will include the proof of Theorem 3. Theorem 4 will be proved in Section 5 where we will present also a number of results regarding the operators $\mathcal{R}_{(\eta)}$ acting on $H^{1}$.
Throughout the paper we shall be using the convention that $C=C(p,\alpha,q,\beta,\dots)$ will denote a positive constant which depends only upon the displayed parameters $p,\alpha,q,\beta\dots$ (which sometimes will be omitted) but not necessarily the same at different occurrences. Furthermore, for two real-valued functions $K_{1},K_{2}$ we write $K_{1}\lesssim K_{2}$, or $K_{1}\gtrsim K_{2}$, if there exists a positive constant $C$ independent of the arguments such that $K_{1}\leq CK_{2}$, respectively $K_{1}\geq CK_{2}$. If we have $K_{1}\lesssim K_{2}$ and $K_{1}\gtrsim K_{2}$ simultaneously, then we say that $K_{1}$ and $K_{2}$ are equivalent and we write $K_{1}\asymp K_{2}$.
2. Preliminary definitions and results
In this section we are going to present a number of spaces and results which will be needed to prove the theorems stated in section 1. Let start fixing some notation. If $f\in\rm{Hol}(\mathbb{D})$, $f(z)=\sum_{k=0}^{\infty}a_{k}z^{k}$ ($z\in\mathbb{D})$, we set
(2.1) $S_{n,m}f(z)\,=\,\sum_{k=n}^{m}a_{k}z^{k},\quad 0\leq n<m<\infty.$
(2.2) $\Delta_{N}f(z)\,=\,S_{N,2N-1}f(z)\,=\,\sum_{k=N}^{2N-1}a_{k}z^{k},\quad 1\leq N<\infty.$
(2.3) $S_{N}f(z)\,=\,\sum_{k=0}^{N}a_{k}z^{k},\quad 0\leq N<\infty.$
2.1. Mean Lipschitz spaces
For $1\leq p\leq\infty$ and $0<\alpha\leq 1$, we let $\Lambda^{p}_{\alpha}$ be the space of those $f\in\rm{Hol}(\mathbb{D})$ having a non-tangential limit at almost every point of $\partial\mathbb{D}$ and so that $\omega_{p}(\cdot,f)$, the integral modulus of continuity of order $p$ of the boundary values $f(e^{i\theta})$ of $f$, satisfies $\omega_{p}(\delta,f)=\operatorname{O}(\delta^{\alpha})$, as $\delta\to 0$. When $p=\infty$ we write $\Lambda_{\alpha}$ instead of $\Lambda^{\infty}_{\alpha}$. This is the usual space of order $\alpha$. More precisely, a function $f\in\rm{Hol}(\mathbb{D})$ belongs to $\Lambda_{\alpha}$ if and only if it has a continuous extension to the closed unit disc and its boundary values satisfy a Lipschitz condition of order $\alpha$.
If $0<\alpha\leq 1$, $1\leq p\leq\infty$, and $f\in\rm{Hol}(\mathbb{D})$, we set
$\beta_{p,\alpha}(r,f)\,=\,(1-r)^{1-\alpha}M_{p}(r,f^{\prime}),\quad 0<r<1.$
Classical results of Hardy and Littlewood (see [10] and [11, Chapter 5]) show that whenever $0<\alpha\leq 1$ and $1\leq p\leq\infty$ we have that $\Lambda^{p}_{\alpha}\subset H^{p}$ and that
(2.4) $\Lambda^{p}_{\alpha}\,=\,\left\{\hbox{$f$ analytic in ${\mathbb{D}}$: $M_{p}(r,f^{\prime})=\operatorname{O}\left((1-r)^{\alpha-1}\right)$, \quad as $r\to 1$}\right\}.$
The corresponding “little oh”  spaces are denoted by $\lambda^{p}_{\alpha}$.
If $1<p<\infty$ and $(1/p)<\alpha\leq 1$, then $\Lambda^{p}_{\alpha}\subset\Lambda_{\alpha-{(1/p)}}$ and, hence, each function in $\Lambda^{p}_{\alpha}$ has a continuous extension to the closed unit disc [10, p. 88]. This is not true for the space $\Lambda^{p}_{1/p}$. This follows easily noticing that the function $f(z)=\log(1-z)$ is an unbounded function which belongs to $\Lambda_{1/p}^{p}$ for all $p\in(1,\infty)$. Bourdon, Shapiro, and Sledd [10] proved that $\Lambda^{p}_{1/p}\subset BMOA$ for all $p\in(1,\infty)$. This inclusion was shown to be sharp in a very strong sense in [9]. Let us remark also that the $\Lambda^{p}_{1/p}$-spaces form a nested scale of spaces
(2.5) $\Lambda^{q}_{1/q}\,\subset\,\Lambda^{p}_{1/p}\quad 1\leq q<p<\infty.$
For $1\leq p\leq\infty$ and $0<\alpha\leq 1$, the space $\Lambda^{p}_{\alpha}$ is a Banach space with the norm $\|\cdot\|_{p,\alpha}$ defined by
$\|f\|_{p,\alpha}\,=\,|f(0)|\,+\,\beta_{p,\alpha}(f),$
where
$\beta_{p,\alpha}(f)\,=\,\sup_{0<r<1}(1-r)^{1-\alpha}M_{p}(r,f^{\prime})\,=\,\sup_{0<r<1}\beta_{p,\alpha}(r,f).$
Let us recall several distinct characterizations of $\Lambda^{p}_{\alpha}$ spaces, (see [10], [11], [18], and [29]).
Theorem A.
Suppose that $1<p<\infty$, $0<\alpha<1$ and $g\in\rm{Hol}({\mathbb{D}})$. The following conditions are equivalent:
  1. (i)
    $g\in\Lambda^{p}_{\alpha}.$
  2. (ii)
    $M_{p}(r,g^{\prime})=\operatorname{O}\left(\frac{1}{(1-r)^{1-\alpha}}\right)$, as $r\to 1^{-}$.
  3. (iii)
    $\|\Delta_{N}g\|_{H^{p}}=\operatorname{O}\left(N^{-\alpha}\right)$, as $N\to\infty$.
  4. (iv)
    $\|\Delta_{N}g^{\prime}\|_{H^{p}}=\operatorname{O}\left(N^{(1-\alpha)}\right)$, as $N\to\infty$.
In fact,
$\beta_{p,\alpha}(f)\,\asymp\,\sup_{N\geq 1}N^{\alpha}\|\Delta_{N}g\|_{H^{p}}\,\asymp\,\sup_{N\geq 1}N^{\alpha-1}\|\Delta_{N}g^{\prime}\|_{H^{p}}.$
Remark 1.
The corresponding results for the little-oh space $\lambda^{p}_{\alpha}$ remain true, and they can be proved following the proofs in the references for Theorem A.
The following result will be needed in our work. Since we have not found any reference to it in the literature, we will include a proof.
Proposition 1.
Suppose that $1<p<\infty$, $0<\alpha\leq 1$, and $f\in\lambda^{p}_{\alpha}$. Then:
(i) Defining, for $0<r<1$, $f_{r}$ by $f_{r}(z)=f(rz)$  ($|z|\leq 1$), we have that
$\beta_{p,\alpha}(f-f_{r})\,\rightarrow 0,\quad\text{as $r\to 1$}.$
(ii) $S_{N}f\,\rightarrow f,\,\,\text{as $N\to\infty$, in the norm of $\Lambda^{p}_{\alpha}$}.$
Proof. Let us start proving (i). Take a non-constant $f\in\lambda^{p}_{\alpha}$. Take $\varepsilon>0$ with $0<\varepsilon<3\beta_{p,\alpha}(f).$ Since $f\in\lambda^{p}_{\alpha}$, there exists $s_{0}\in(0,1)$ such that
(2.6) $(1-s)^{1-\alpha}M_{p}(s,f^{\prime})\,<\frac{\varepsilon}{3},\quad s_{0}\leq s<1.$
For $r,s\in(0,1)$, we have
(2.7) $(1-s)^{1-\alpha}M_{p}(s,f^{\prime}-f^{\prime}_{r})\,=\,(1-s)^{1-\alpha}\left[\frac{1}{2\pi}\int_{0}^{2\pi}\left|f^{\prime}(se^{i\theta})-rf^{\prime}(rse^{i\theta})\right|^{p}\,d\theta\right]^{1/p}$
$=\,$ $(1-s)^{1-\alpha}\left[\frac{1}{2\pi}\int_{0}^{2\pi}\left|f^{\prime}(se^{i\theta})-rf^{\prime}(se^{i\theta})+rf^{\prime}(se^{i\theta})-rf^{\prime}(rse^{i\theta})\right|^{p}\,d\theta\right]^{1/p}$
$\leq\,$ $(1-r)(1-s)^{1-\alpha}M_{p}(s,f^{\prime})\,+\,r(1-s)^{1-\alpha}\left[\frac{1}{2\pi}\int_{0}^{2\pi}\left|f^{\prime}(se^{i\theta})-f^{\prime}(rse^{i\theta})\right|^{p}\,d\theta\right]^{1/p}.$
Using this, (2.6), and the fact that $M_{p}(rs,f^{\prime})\leq M_{p}(s,f^{\prime})$ ($0<s,r<1$), we obtain
$(1-s)^{1-\alpha}M_{p}(s,f^{\prime}-f^{\prime}_{r})\,\leq\,$ $(1-s)^{1-\alpha}M_{p}(s,f^{\prime})+\,(1-s)^{1-\alpha}\left[M_{p}(s,f^{\prime})\,+\,M_{p}(rs,f^{\prime})\right]$
(2.8) $\leq\,$ $3(1-s)^{1-\alpha}M_{p}(s,f^{\prime})$
$<\,$ $\varepsilon,\qquad s_{0}<s<1.$
Since $f^{\prime}$ is uniformly continuous in the closed disc $\{z\in\mathbb{D}:|z|\leq s_{0}\}$, there exists $\delta\in(0,1)$ such that
(2.9) $|f^{\prime}(z)-f^{\prime}(w)|<\frac{2\varepsilon}{3},\,\,\text{if $|z|\leq s_{0},\,|w|\leq s_{0}$,\,\,\,and $|z-w|\leq\delta$}.$
Take $r_{0}\in(0,1)$ with
$r_{0}>\max\left(1-\delta,1-\frac{\varepsilon}{3\beta_{p,\alpha}(f)}\right).$
Using (2.7) and (2.9), we see that, if $0<s\leq s_{0}$ and $r_{0}<r<1$, then
$(1-s)^{1-\alpha}M_{p}(s,f^{\prime}-f^{\prime}_{r})\leq\,(1-r)\beta_{p,\alpha}(f)\,+\,(1-s)^{1-\alpha}\frac{2\varepsilon}{3}\leq\frac{\varepsilon}{3\beta_{p,\alpha}(f)}\beta_{p,\alpha}(f)+\,\frac{2\varepsilon}{3}=\,\varepsilon.$
This and (2.1) give that $\beta_{p,\alpha}(f-f_{r})<\varepsilon$, if $r_{0}<r<1$. Hence, (i) is proved.
In order to prove (ii), we shall start proving the following weaker result.
Lemma 1.
Suppose that $1<p<\infty$, $0<\alpha\leq 1$, and $g\in\lambda^{p}_{\alpha}$. Then $g$ is the limit in $\Lambda^{p}_{\alpha}$ of a sequence of polynomials.
Proof. Take $g\in\lambda^{p}_{\alpha}$. Set $r_{n}=1-\frac{1}{n}$ ($n=2,3,\dots$) and consider the sequence of functions $\{g_{r_{n}}\}_{n=2}^{\infty}$. For every $n\geq 2$, $g_{r_{n}}$ is holomorphic in $\{|z|\leq\frac{1}{r_{n}}\}$ and, hence, there exists a polynomial $P_{n}$ with $P_{n}(0)=g(0)$ such that
$|g^{\prime}_{r_{n}}(z)\,-\,P^{\prime}_{n}(z)|\leq\frac{1}{n},\quad|z|\leq 1.$
Clearly, this implies that
$\beta_{p,\alpha}(g_{r_{n}}-P_{n})\leq\frac{1}{n},\quad n\geq 2.$
Using this and (i) we see that
$\beta_{p,\alpha}(g-P_{n})\,\leq\beta_{p,\alpha}(g-g_{r_{n}})\,+\,\beta_{p,\alpha}(g_{r_{n}}-P_{n})\,\to 0,\quad\text{as $n\to\infty$}.$
Thus, $P_{n}\,\to g$, as $n\to\infty$, in the norm of $\Lambda^{p}_{\alpha}$. $\Box$
Bearing in mind Lemma 1 and Proposition  1 of [42], (ii) will follow if we prove that there exists a positive constant $C$ such that
(2.10) $\beta_{p,\alpha}(S_{n}f)\leq C\beta_{p,\alpha}(f),\,\text{ for all $f\in\Lambda^{p}_{\alpha}$ and all $n\in\mathbb{N}$.}$
Since $1<p<\infty$, using the Riesz projection theorem, it follows that the operators $S_{n}$ are uniformly bounded on $H^{p}$, that is, there exists $A>0$ such that
(2.11) $\|S_{n}f\|_{H^{p}}\leq A\|f\|_{H^{p}},\quad(n\in\mathbb{N},\,\,f\in H^{p}).$
Take now $f\in\Lambda^{p}_{\alpha}$, $f(z)\,=\,\sum_{k=0}^{\infty}a_{k}z^{k}$ ($z\in\mathbb{D}$). Notice that for $n\in\mathbb{N}$ we have
$\left(S_{n}f\right)^{\prime}(z)\,=\,\sum_{k=0}^{n-1}(k+1)a_{k+1}z^{k}\,=\,S_{n-1}(f^{\prime})(z),\quad z\in\mathbb{D}.$
Then, using (2.11), we see that
$\beta_{p,\alpha}(S_{n}f)\,=\,\sup_{0<r<1}(1-r)^{1-\alpha}M_{p}(r,(S_{n}f)^{\prime})$
$=\,$ $\sup_{0<r<1}(1-r)^{1-\alpha}M_{p}(r,S_{n-1}(f^{\prime}))\,\leq A\sup_{0<r<1}(1-r)^{1-\alpha}M_{p}(r,f^{\prime})\,=\,A\beta_{p,\alpha}(f).$
Hence, (2.10) holds with $C=A$. This finishes the proof. $\Box$
2.2. Convolution of analytic functions
If $f$ and $g$ are two analytic functions in the unit disc,
$f(z)=\sum_{n=0}^{\infty}a_{n}z^{n},\quad g(z)=\sum_{n=0}^{\infty}b_{n}z^{n},\quad z\in\mathbb{D},$
the convolution $f\star g$ of $f$ and $g$ is defined by
(2.12) $f\star g(z)\,=\,\sum_{n=0}^{\infty}a_{n}b_{n}z^{n},\quad z\in\mathbb{D}.$
We have
(2.13) $f\star g(\rho re^{i\theta})\,=\,\frac{1}{2\pi}\int_{0}^{2\pi}f(\rho e^{it})g(re^{i(\theta-t)})\,dt,\quad 0<\rho<1,\,\,0<r<1,\,\,\theta\in\mathbb{R}.$
It is well known that the convolution of a function in $H^{1}$ and another one in $H^{p}$ ($p\geq 1$) lies in $H^{p}$ and that
$\|f\star g\|_{H^{p}}\leq\|f\|_{H^{1}}\|g\|_{H^{p}},\quad p\geq 1,f\in H^{1},g\in H^{p}.$
Actually, (2.13) implies that if $f\in H^{1}$ and $g\in\rm{Hol}(\mathbb{D})$, then
(2.14) $M_{p}(r,f\star g)\,\leq\|f\|_{H^{1}}M_{p}(r,g),\quad 0<r<1.$
Also, taking $\rho=r$ in (2.13), we obtain
(2.15) $M_{p}(r^{2},f\star g)\,\leq M_{1}(r,f)M_{p}(r,g),\quad f,g\in\rm{Hol}(\mathbb{D}),\,\,0<r<1.$
We shall use also the following elementary result.
Lemma 2.
Suppose that $\alpha>-1$, $1\leq p<\infty$, $f\in H^{p}$, and $g\in A^{p}_{\alpha}$. Then
$\|f\star g\|_{A^{p}_{\alpha}}\,\lesssim\|f\|_{H^{p}}\|g\|_{A^{p}_{\alpha}}.$
Proof.
$\|f\ast g\|_{A^{p}_{\alpha}}^{p}$ $\asymp\int_{0}^{1}(1-r)^{\alpha}M_{p}^{p}(r,f\ast g)\,dr\,\lesssim\int_{0}^{1}(1-r)^{\alpha}\|f\|_{H^{1}}^{p}M_{p}^{p}(r,g)\,dr$
$\lesssim\|f\|_{H^{p}}^{p}\int_{0}^{1}(1-r)^{\alpha}M_{p}^{p}(r,g)\,dr\,\asymp\|f\|_{H^{p}}^{p}\|g\|_{A^{p}_{\alpha}}^{p}.$
$\Box$
2.3. “Lipschitz” Cesàro bases
Induced by the work of Pavlović [34, pp.  140-143] who constructed “smooth”  Cesàro bases, next we are going to construct “Lipschitz” Cesàro bases.
Let $\Psi:\mathbb{R}\rightarrow\mathbb{C}$ be a Lipschitz function with compact support. For simplicity, we shall assume that the support of $\Psi$ is contained in $[0,4]$. We set
$\mathcal{L}(\psi)\,=\,\sup_{s\neq t}\frac{|\Psi(t)-\Psi(s)|}{|t-s|}.$
For $n\in\mathbb{N}$ we define
$W_{n}^{\Psi}(e^{it})\,=\,\sum_{k=-\infty}^{\infty}\Psi\left(\frac{k}{n}\right)e^{ik\theta}\,=\,\sum_{k=0}^{4n}\Psi\left(\frac{k}{n}\right)e^{ik\theta}.$
Arguing as in the proof of Lemma 5.2 in p. 141 of [34] with $N=2$, we have
(2.16) $\,(1-e^{i\theta})^{2}W_{n}^{\Psi}(e^{i\theta})\,=\,\sum_{k=0}^{4n}\Psi\left(\frac{k}{n}\right)(1-e^{i\theta})^{2}e^{ik\theta}\,=\,\sum_{k=0}^{4n}\Psi\left(\frac{k}{n}\right)\sum_{m=0}^{2}\binom{2}{m}(-1)^{m}e^{i(m+k)\theta}$
$=\,$ $\sum_{m=0}^{2}(-1)^{m}\binom{2}{m}\sum_{k=0}^{4n}\Psi\left(\frac{k}{n}\right)e^{i(m+k)}\theta\,=\,\sum_{m=0}^{2}(-1)^{m}\binom{2}{m}\sum_{j=0}^{4n+2}\Psi\left(\frac{j-m}{n}\right)e^{ij\theta}$
$=\,$ $\sum_{j=0}^{4n+2}\left[\sum_{m=0}^{2}(-1)^{m}\binom{2}{m}\Psi\left(\frac{j-m}{n}\right)\right]e^{ij\theta}.$
Now,
$\left|\sum_{m=0}^{2}(-1)^{m}\binom{2}{m}\Psi\left(\frac{j-m}{n}\right)\right|=$ $\left|\Psi\left(\frac{j}{n}\right)-\Psi\left(\frac{j-1}{n}\right)-\Psi\left(\frac{j-1}{n}\right)+\Psi\left(\frac{j-2}{n}\right)\right|\leq\frac{2\mathcal{L}(\psi)}{n}.$
Using this in (2.16), we obtain
(2.17) $\left|(1-e^{i\theta})^{2}W_{n}^{\Psi}(e^{i\theta})\right|\leq\frac{2\mathcal{L}(\Psi)}{n}\sum_{j=0}^{4n+2}|e^{ij\theta}|\leq\frac{2\mathcal{L}(\Psi)}{n}(4n+3)\leq 14\mathcal{L}(\Psi).$
3. Rhaly operators acting on Hardy spaces
Let us start proving that, for $1<p<\infty$, the condition $F_{(\eta)}\in\Lambda^{p}_{1/p}$ (respectively, $F_{(\eta)}\in\lambda^{p}_{1/p}$) is a necessary condition for the membership of $\mathcal{R}_{(\eta)}$ in $\mathcal{B}(H^{p})$ (respectively, in $\mathcal{K}(H^{p}))$.
Proof. Assume that $1<p<\infty$ and $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$. As observed in section 1 this implies that $F_{(\eta)}$ is a well defined analytic function in $\mathbb{D}$. In fact, since $F_{(\eta)}=\mathcal{R}_{(\eta)}(1)$, $F_{(\eta)}\in H^{p}$.
For $N=2,3,\dots$, set $a_{N}=1-\frac{1}{N}$ and
(3.1) $f_{N}(z)\,=\,\frac{1}{N^{2-\frac{1}{p}}}\frac{a_{N}z}{(1-a_{N}z)^{2}}\,=\,\frac{1}{N^{2-\frac{1}{p}}}\sum_{n=1}^{\infty}na_{N}^{n}z^{n},\quad z\in\mathbb{D}.$
We have that $f_{N}\in H^{p}$ for all $N$ and
(3.2) $\|f_{N}\|_{H^{p}}\lesssim 1.$
Since $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$,
(3.3) $\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}\lesssim\|f_{N}\|_{H^{p}}.$
Now,
$\mathcal{R}_{(\eta)}(f_{N})(z)\,=\,\sum_{k=1}^{\infty}\eta_{k}\left(\frac{1}{N^{2-\frac{1}{p}}}\sum_{n=1}^{k}na_{N}^{n}\right)z^{k}\,=\,\sum_{k=1}^{\infty}\eta_{k}k\left(\frac{1}{kN^{2-\frac{1}{p}}}\sum_{n=1}^{k}na_{N}^{n}\right)z^{k}.$
For $k,N\in\mathbb{N}$, set
(3.4) $\alpha_{k,N}\,=\,\frac{1}{kN^{2-\frac{1}{p}}}\sum_{n=1}^{k}na_{N}^{n},\quad\beta_{k,N}\,=\,\frac{1}{\alpha_{k,N}}.$
Then
(3.5) $\mathcal{R}_{(\eta)}(f_{N})(z)\,=\,\sum_{k=1}^{\infty}\eta_{k}k\alpha_{k,N}z^{k},\quad z\in\mathbb{D}.$
Let us point out some properties of the $\alpha_{k,N}$’s and the $\beta_{k,N}$’s.
(3.6) $0\,\leq\,\alpha_{k,N}\,\leq\,\frac{1}{N^{2-\frac{1}{p}}}\frac{1}{k}\sum_{n=1}^{k}n\,=\,\frac{1}{N^{2-\frac{1}{p}}}\frac{k+1}{2}.$
(3.7) $\alpha_{k,N}\geq\frac{1}{N^{2-\frac{1}{p}}}a_{N}^{k}\frac{1}{k}\sum_{n=1}^{k}n\,=\,\frac{a_{N}^{k}}{N^{2-\frac{1}{p}}}\frac{k+1}{2}.$
Using (3.6) and (3.7), we see that there exist two positive constants $C_{1}$ and $C_{2}$ such that
(3.8) $\frac{C_{1}}{N^{1-\frac{1}{p}}}\,\leq\,\alpha_{k,N}\,\leq\,\frac{C_{2}}{N^{1-\frac{1}{p}}},\quad N=,2,3,\dots,\quad N\leq k\leq 2N.$
We have also, for all $k,N$,
$|\alpha_{k+1,N}-\alpha_{k,N}|\,=\,$ $\frac{1}{N^{2-\frac{1}{p}}}\left|\frac{1}{k+1}\sum_{n=1}^{k+1}na_{N}^{n}\,-\,\frac{1}{k}\sum_{n=1}^{k}a_{N}^{n}\right|$
[…]
$\leq\,$ $\frac{3}{2N^{2-\frac{1}{p}}}.$
Using this, (3.6), (3.7), and (3.8), we deduce that
(3.9) $|\beta_{k+1,N}-\beta_{k,N}|\,=\,\frac{|\alpha_{k+1,N}-\alpha_{k,N}|}{|\alpha_{k+1,N}\alpha_{k,N}|}\,\lesssim N^{-\frac{1}{p}},\quad N\leq k\leq 2N,$
and then, we obtain,
(3.10) $\left|\frac{\beta_{k+1,N}-\beta_{k,N}}{\frac{k+1}{N}-\frac{k}{N}}\right|\,=\,N\left|\beta_{k+1,N}-\beta_{k,N}\right|\,\lesssim N^{1-\frac{1}{p}},\quad N\leq k\leq 2N.$
For each $N$, let $\Psi_{N}:\mathbb{R}\rightarrow\mathbb{R}$ be defined as follows:
  • $\Psi_{N}(x)=0$, if $x\leq 0$.
  • $\Psi_{N}(x)=0$, if $x\geq 4$.
  • Between 0 and $4$, the graph of $\Psi_{N}$ is the polygonal whose vertices are the points
    $(0,0),\,\left(\frac{N}{N},\beta_{N,N}\right),\,\left(\frac{N+1}{N},\beta_{N+1,N}\right),\,\left(\frac{N+2}{N},\beta_{N+2,N}\right),\,\dots,\left(\frac{2N}{N},\beta_{2N,N}\right),\,(4,0).$
Using (3.10), we see that $\Psi_{N}$ is a Lipschitz function and that
(3.11) $\mathcal{L}(\Psi_{N})\,\lesssim N^{1-\frac{1}{p}}.$
Taking $\Psi=\Psi_{N}$ and $n=N$ in section 2.3, and using (2.17), we obtain
(3.12) $\left|\sum_{k=0}^{4N}\Psi_{N}\left(\frac{k}{N}\right)e^{ik\theta}\right|\,\lesssim N^{1-\frac{1}{p}}\frac{1}{|1-e^{i\theta}|^{2}},\quad 0<\theta<2\pi.$
For each $N$ define $H_{N}$ by
$H_{N}(z)\,=\,\sum_{k=0}^{4N}\Psi_{N}\left(\frac{k}{N}\right)z^{k},\quad z\in\mathbb{D}.$
Using (3.12), we deduce that
(3.13) $\|H_{N}\|_{H^{p}}\lesssim N^{1-\frac{1}{p}}.$
Using (2.14), the fact that the $\|\cdot\|_{H^{p}}$ norm increases with $p$, and (3.13), it follows that
(3.14) $\|H_{N}\star\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}\leq$ $\|H_{N}\|_{H^{1}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}\lesssim\|H_{N}\|_{H^{p}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}\lesssim N^{1-\frac{1}{p}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}.$
Set
$G_{(\eta)}(z)=zF^{\prime}_{(\eta)}(z)=\sum_{k=1}^{\infty}k\eta_{k}z^{k},\quad z\in\mathbb{D}.$
Bearing in mind the definitions of $\Psi_{N}$ and $H_{N}$, we see that
$\Delta_{N}(G_{(\eta)})(z)\,=\,\sum_{k=N}^{2N-1}k\eta_{k}z^{k}$
$=\,$ $H_{N}\star\mathcal{R}_{(\eta)}(f_{N})(z)\,-S_{N-1}(H_{N}\star\mathcal{R}_{(\eta)})(f_{N}))(z)\,-\,S_{2N,4N+1}(H_{N}\star\mathcal{R}_{(\eta)})(f_{N}))(z)$
$=\,$ $H_{N}\star\mathcal{R}_{(\eta)}(f_{N})(z)\,-S_{N-1}\left(H_{N}\star\mathcal{R}_{(\eta)}(f_{N})\right)(z)$
$-S_{4N+1}\left(H_{N}\star\mathcal{R}_{(\eta)}(f_{N})\right)(z)+S_{2N-1}\left(H_{N}\star\mathcal{R}_{(\eta)}(f_{N})\right))(z).$
Then, using (3.14) and the Riesz projection theorem, we deduce that
$\|\Delta_{N}(G_{(\eta)})\|_{H^{p}}\lesssim N^{1-\frac{1}{p}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}.$
Clearly, this implies that
(3.15) $\|\Delta_{N}(F^{\prime}_{(\eta)})\|_{H^{p}}\,\lesssim N^{1-\frac{1}{p}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}.$
Using (3.3), (3.2), (3.15), and Theorem A, we obtain that $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
Suppose now that $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$. Since $f_{N}\to 0$, as $N\to\infty$, uniformly in compact subsets of $\mathbb{D}$, using Lemma 3. 7 of [40], we see that $\mathcal{R}_{(\eta)}(f_{N})\to 0$, as $N\to\infty$, in $H^{p}$. This and (3.15) give that
$\|\Delta_{N}(F^{\prime}_{(\eta)})\|_{H^{p}}\,=\,\operatorname{o}\left(N^{1-\frac{1}{p}}\right),\quad\text{as $N\to\infty$}.$
This gives that $F_{(\eta)}\in\lambda^{p}_{1/p}$$\Box$
In order to prove the sufficiency statements in Theorem 1 and Theorem 2 we will work with the spaces $\mathcal{D}^{p}_{p-1}$ ($1\leq p<\infty$) and with the spaces, which we will call $\mathcal{X}_{q,p}$, consisting of those $f\in\rm{Hol}(\mathbb{D})$ whose derivative belongs to the mixed norm space $H^{q,p,\alpha}$ with $\alpha=1+\frac{1}{p}-\frac{1}{q}$ ($1\leq q\leq p<\infty$). Thus,
$\mathcal{X}_{q,p}=\left\{f\in\rm{Hol}(\mathbb{D}):\|f\|_{\mathcal{X}_{q,p}}\stackrel{{\scriptstyle\text{def}}}{{=}}\left(|f(0)|^{p}\,+\,\int_{0}^{1}(1-r)^{p(1-\frac{1}{q})}M_{q}^{p}(r,f^{\prime})\,dr\right)^{1/p}<\infty\right\}.$
Notice that
(3.16) $\mathcal{X}_{p,p}\,=\,\mathcal{D}^{p}_{p-1},\quad 1\leq p<\infty.$
It is well known (see, e. g., [27, 28, 41]) that
(3.17) $\mathcal{D}^{p}_{p-1}\subset H^{p},\quad 0<p\leq 2,$
and
(3.18) $H^{p}\subset\mathcal{D}^{p}_{p-1},\quad 2<p<\infty,$
the inclusions being continuous.
A result of Flett [13, Theorem 1] (see also Corollary 3. 1 of [29]) asserts that
(3.19) If $1<q<p$ then $\mathcal{X}_{q,p}$ is continuously embedded in $H^{p}$.
Bearing in mind (3.17) and (3.19), the sufficiency statements in Theorem 1 and Theorem 2 will follow readily from the following result.
Theorem 5.
Let $(\eta)\,=\,\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers.
(i) If  $1\leq p<\infty$ and $F_{(\eta)}\in\Lambda^{p}_{1/p}$, then $\mathcal{R}_{(\eta)}$ is a bounded operator from $H^{p}$ into $\mathcal{D}^{p}_{p-1}$.
(ii) If  $1<p<\infty$ and $F_{(\eta)}\in\lambda^{p}_{1/p}$, then $\mathcal{R}_{(\eta)}$ is a compact operator from $H^{p}$ into $\mathcal{D}^{p}_{p-1}$.
(iii) If  $1<q<p<\infty$ and $F_{(\eta)}\in\lambda^{q}_{1/q}$, them $\mathcal{R}_{(\eta)}$ is a compact operator from $H^{p}$ into $\mathcal{X}_{q,p}$.
Our proof of this result is inspired in the work of M. Nowak [32] who proved that the Cesàro operator is bounded from $H^{p}$ into $\mathcal{D}^{p}_{p-1}$ ($0<p<\infty$).
In order to simplify the notation, we shall let $S$ denote the shift operator on $\rm{Hol}(\mathbb{D})$, that is,
(3.20) $S[f](z)\,=\,zf(z),\quad z\in\mathbb{D},\,\,\,f\in\rm{Hol}(\mathbb{D}).$
Also, let us notice that, for $h\in\rm{Hol}(\mathbb{D})$, we have
(3.21) $\|h\|_{\mathcal{X}_{q,p}^{p}}\,\asymp\,$ $|h(0)|^{p}\,+\,\int_{0}^{1}(1-r)^{p\left(1-\frac{1}{q}\right)}M_{q}^{p}(r,h^{\prime})\,dr$
$=\,$ $|h(0)|^{p}\,+\,\int_{0}^{1}2r(1-r^{2})^{p\left(1-\frac{1}{q}\right)}M_{q}^{p}(r^{2},h^{\prime})\,dr$
$\lesssim\,$ $|h(0)|^{p}\,+\,\int_{0}^{1}(1-r)^{p\left(1-\frac{1}{q}\right)}M_{q}^{p}(r^{2},h^{\prime})\,dr.$
Proof ofTheorem 5. Suppose that $1\leq q\leq p<\infty$ and $F_{(\eta)}\in\Lambda^{q}_{1/q}$.
Take $f\in H^{p}$, $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ ($z\in\mathbb{D}$) and set
(3.22) $g(z)\,=\,\frac{f(z)}{1-z}\,=\,\sum_{n=0}^{\infty}\left(\sum_{k=0}^{n}a_{k}\right)z^{n},\quad z\in\mathbb{D}.$
We have,
$S\left[\mathcal{R}_{(\eta)}(f)\right](z)\,=\,\sum_{n=0}^{\infty}(n+1)\eta_{n}\left(\frac{1}{n+1}\sum_{k=0}^{n}a_{k}\right)z^{n+1},\quad z\in\mathbb{D},$
and then
(3.23) $S\left[\mathcal{R}_{(\eta)}(f)\right]^{\prime}(z)\,=\,\sum_{n=0}^{\infty}(n+1)\eta_{n}\left(\sum_{k=0}^{n}a_{k}\right)z^{n}\,=\,S\left[F_{(\eta)}\right]^{\prime}\star g(z),\quad z\in\mathbb{D}.$
Using (2.15) and the continuity of the shift operator, we obtain, for $1/2<r<1$,
(3.24) $M_{q}\left(r^{2},\left[\mathcal{R}_{(\eta)}(f)\right]^{\prime}\right)$ $\,\lesssim M_{q}\left(r,S\left[\mathcal{R}_{(\eta)}(f)\right]^{\prime}\right)$
$\,\lesssim M_{q}\left(r,S[F_{(\eta)}]^{\prime}\right)M_{1}(r,g)$
$\,\lesssim M_{q}\left(r,F_{\eta}^{\prime}\right)M_{1}(r,g)$
$\lesssim(1-r)^{\frac{1}{q}-1}\beta_{q,\frac{1}{q}}\left(r,F_{(\eta)}\right)M_{1}(r,g).$
Take $\gamma>1$ and let $\sigma$ be the exponent conjugate to $\gamma p$, that is, $\frac{1}{\gamma p}+\frac{1}{\sigma}=1$. Using Hölder‘s inequality, we obtain
$M_{1}(r,g)=\frac{1}{2\pi}\int_{0}^{2\pi}\left|\frac{f(re^{i\theta}}{1-re^{i\theta}}\right|d\theta\,\leq M_{\gamma p}(r,f)\left(\frac{1}{2\pi}\int_{0}^{2\pi}\frac{d\theta}{|1-re^{i\theta}|^{\sigma}}\right)^{1/\sigma}\lesssim\frac{1}{(1-r)^{\frac{1}{\gamma p}}}M_{\gamma p}(r,f).$
This and (3.24) give
$M_{q}\left(r^{2},\left[\mathcal{R}_{(\eta)}(f)\right]^{\prime}\right)$ $\,\lesssim(1-r)^{\frac{1}{q}-1}\beta_{q,\frac{1}{q}}\left(r,F_{(\eta)}\right)\frac{1}{(1-r)^{\frac{1}{\gamma p}}}M_{\gamma p}(r,f).$
This and (3.21) imply that
(3.25) $\|\mathcal{R}_{(\eta)}(f)\|_{\mathcal{X}_{q,p}}^{p}\,\lesssim$ $|\eta_{0}|^{p}|a_{0}|^{p}\,+\int_{0}^{1}(1-r)^{p\left(1-\frac{1}{q}\right)}M_{q}^{p}\left(r^{2},\left[\mathcal{R}_{(\eta)}(f)\right]^{\prime}\right)\,dr$
$\lesssim$ $|\eta_{0}|^{p}\|f\|_{H^{p}}^{p}\,+\,\int_{0}^{1}\beta^{p}_{q,\frac{1}{q}}\left(r,F_{(\eta)}\right)\frac{1}{(1-r)^{\frac{1}{\gamma}}}M_{\gamma p}^{p}(r,f)dr.$
Since $F_{(\eta)}\in\Lambda^{q}_{1/q}$,   $\sup_{0<r<1}\beta_{q}\left(r,F_{(\eta)}\right)<\infty$. Then, using Theorem 5. 11 of [11], it follows that
(3.26) $\|\mathcal{R}_{(\eta)}(f)\|_{\mathcal{X}_{q,p}}^{p}\lesssim\|f\|_{H^{p}}^{p}.$
Thus we have proved that $\mathcal{R}_{(\eta)}$ is bounded from $H^{p}$ into $\mathcal{X}_{q,p}$.
Part (i) of Theorem 5 follows taking $q=p$ and using (3.16).
Parts (ii) and (iii) of Theorem 5 follow from the following result.
Proposition 2.
Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers. If $1<q\leq p<\infty$ and $f\in\lambda^{q}_{1/q}$, then $\mathcal{R}_{(\eta)}$ is compact form $H^{p}$ into $\mathcal{X}_{q,p}$.
Proof. Suppose that $1<q\leq p<\infty$ and $f\in\lambda^{p}_{1/p}$. We already know that $\mathcal{R}_{(\eta)}$ is bounded from $H^{p}$ into $\mathcal{X}_{q,p}$. For $N=2,3,4,\dots$ let $\left(\eta^{(N)}\right)\,=\,\{\eta^{(N)}_{n}\}_{n=0}^{\infty}$ be the sequence defined by $\eta^{(N)}_{n}=0$, if $0\leq n\leq N$, and $\eta^{(N)}_{n}=\eta_{n}$, if $n\geq N+1$. Set also $\mathcal{R}_{N}:\,\rm{Hol}(\mathbb{D})\rightarrow\rm{Hol}(\mathbb{D})$ be the operator defined by
(3.27) $\mathcal{R}_{N}(f)(z)\,=\,\sum_{n=0}^{N}\eta_{n}\left(\sum_{k=0}^{n}a_{k}\right)z^{n},\quad z\in\mathbb{D},\,\,\,f\in\rm{Hol}(\mathbb{D}).$
The operators $\mathcal{R}_{N}$ are finite rank operators from $H^{p}$ into $\mathcal{X}_{q,p}$. Then it suffices to prove that
(3.28) $\mathcal{R}_{N}\,\rightarrow\,\mathcal{R}_{(\eta)},\quad\text{as $N\to\infty$,\, in the operator norm}.$
Notice that
$\mathcal{R}_{(\eta)}\,-\,\mathcal{R}_{N}\,=\,\mathcal{R}_{(\eta^{(N)})}\,\,\,\text{and}\,\,\,F_{(\eta^{(N)})}\,=\,F_{(\eta)}\,-\,S_{N}F_{(\eta)}.$
Take $\gamma>1$. Then, using (3.25) with $(\eta^{(N)})$ in the place of $(\eta)$, and Theorem 5. 11 of [11] as above, we see that if $f\in H^{p}$ then
$\|\left(\mathcal{R}_{(\eta)}\,-\,\mathcal{R}_{N}\right)(f)\|_{\mathcal{X}_{q,p}}^{p}\,\lesssim\,$ $\int_{0}^{1}\beta_{q,\frac{1}{q}}^{p}\left(r,F_{(\eta)}\,-\,S_{N}F_{(\eta)}\right)\frac{1}{(1-r)^{\frac{1}{\gamma}}}M_{\gamma p}^{p}(r,f)dr$
$\,\lesssim$ $\,\beta_{q,\frac{1}{q}}^{q}\left(F_{(\eta)}-S_{n}F_{(\eta)}\right)\int_{0}^{1}\frac{1}{(1-r)^{\frac{1}{\gamma}}}M_{\gamma p}^{p}(r,f)dr$
$\,\lesssim$ $\,\beta_{q,\frac{1}{q}}^{q}\left(F_{(\eta)}-S_{n}F_{(\eta)}\right)\|f\|_{H^{p}}^{p}.$
Using Proposition 1, (3.28) follows as desired. $\Box$ $\Box$
4. Rhaly operators acting on Bergman spaces
Theorem 3 follows from the following one.
Theorem 6.
Let $(\eta)\,=\,\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers.
(i) If $\alpha>-1$, $1\leq p<\infty$, and $F_{(\eta)}\in\Lambda^{p}_{1/p}$ then $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p}_{\alpha})$.
(ii) If $\alpha>-1$, $1<p<\infty$, and $F_{(\eta)}\in\lambda^{p}_{1/p}$ then $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p}_{\alpha})$.
(iii) If $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p}_{\alpha})$ then $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
(iv) If $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p}_{\alpha})$ then $F_{(\eta)}\in\lambda^{p}_{1/p}$.
Proof. Let us start proving (i) and (ii). So, suppose that $1\leq p<\infty$, $\alpha>-1$, and $F_{(\eta)}\in\Lambda^{p}_{1/p}$. Take $f\in A^{p}_{\alpha}$, $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ ($z\in\mathbb{D}$). Just as in the proof of Theorem 5, set $g(z)=\frac{f(z)}{1-z}$ ($z\in\mathbb{D}$). Notice that using Hölder’s inequality we have
$M_{1}(r,g)\,\lesssim\frac{1}{(1-r)^{1/p}}M_{p}(r,f).$
Using these facts, (3.23), Theorem 4. 28 of [43], the analogue of (3.21) for the space $A^{p}_{p+\alpha}$, and (2.15), we obtain,
(4.1) $\left\|\mathcal{R}_{(\eta)}(f)\right\|_{A^{p}_{\alpha}}^{p}$ $\,\asymp\,\left|\mathcal{R}_{(\eta)}(f)(0)\right|^{p}\,+\,\int_{\mathbb{D}}(1-|z|)^{p+\alpha}\left|\left[\mathcal{R}_{(\eta)}(f)\right]^{\prime}(z)\right|^{p}dA(z)$
$\lesssim|\eta_{0}|^{p}|a_{0}|^{p}\,+\,\int_{0}^{1}(1-r)^{p+\alpha}M_{p}^{p}\left(r^{2},\left[R_{(\eta)}(f)\right]^{\prime}\right)\,dr$
$\lesssim|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{p+\alpha}M_{p}^{p}\left(r^{2},F_{(\eta)}^{\prime}\star g\right)\,dr$
$\lesssim|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{p+\alpha}M_{p}^{p}\left(r,F_{(\eta)}^{\prime}\right)M_{1}^{p}(r,g)\,dr$
$\lesssim|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{p+\alpha}(1-r)^{1-p}\beta_{p,\frac{1}{p}}^{p}\left(r,F_{(\eta)}\right)\frac{M_{p}^{p}(r,f)}{1-r}\,dr$
$=|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{\alpha}\beta_{p,\frac{1}{p}}^{p}(r,F_{(\eta)})M_{p}^{p}(r,f)\,dr.$
Then, using the fact that $F_{(\eta)}\in\Lambda^{p}_{1/p}$, we obtain that
$\left\|\mathcal{R}_{(\eta)}(f)\right\|_{A^{p}_{\alpha}}^{p}\,\lesssim\,|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{\alpha}M_{p}^{p}(r,f)\,dr\,\lesssim\|f\|_{A^{p}_{\alpha}}^{p}.$
Hence, we have proved (i).
Suppose now that $1<p<\infty$ and $F_{(\eta)}\in\lambda^{p}_{1/p}$. For $N=2,3,4,\dots$ let $\mathcal{R}_{N}$ be defined as in (3.27). Using (4.1) and arguing as in the proof of Proposition 2, we deduce, for $f\in A^{p}_{\alpha}$,
$\left\|\left(\mathcal{R}_{(\eta)}\,-\,\mathcal{R}_{N}\right)(f)\right\|_{A^{p}_{\alpha}}^{p}\lesssim$ $\int_{0}^{1}\beta^{p}_{p,\frac{1}{p}}\left(r,F_{(\eta)}\,-\,S_{N}F_{(\eta)}\right)(1-r)^{\alpha}M_{p}^{p}(r,f)^{p}\,dr$
$\lesssim$ $\,\beta^{p}_{p,\frac{1}{p}}\left(F_{(\eta)}\,-\,S_{N}F_{(\eta)}\right)\|f\|_{A^{p}_{\alpha}}^{p}.$
Using Proposition 1 we deduce that $\{\mathcal{R}_{(\eta)}-\mathcal{R}_{N}\}\,\rightarrow\,0$, as $N\to\infty$, in the norm of $\mathcal{B}(A^{p}_{\alpha})$. Since the operators $\mathcal{R}_{N}:A^{p}_{\alpha}\rightarrow A^{p}_{\alpha}$ are finite rank operators, it follows that $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p}_{\alpha})$. Thus, (ii) is proved.
The proof of parts (iii) and (iv) is closely related to that of the necessity statements in Theorems 1 and 2.
Suppose that $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p}_{\alpha})$.
For $N=2,3,4,\dots$, set $a_{N}=1-\frac{1}{N}$ and
$g_{N}(z)\,=\,\frac{1}{N^{2-\frac{\alpha+2}{p}}}\frac{a_{N}z}{(1-a_{N}z)^{2}}\,=\,N^{\frac{\alpha+1}{p}}f_{N}(z),\quad z\in\mathbb{D},$
where $f_{N}$ is defined in (3.1). Since $\alpha<2p-2$, Lemma 3. 10 of [43] shows that $g_{N}\in A^{p}_{\alpha}$ for all $N$ and $\|g_{N}\|_{A^{p}_{\alpha}}\asymp 1$. For $k,N\in\mathbb{N}$, set
(4.2) $\gamma_{k,N}\,=\,N^{\frac{1+\alpha}{p}}\alpha_{k,N},\quad\delta_{k,N}\,=\,\frac{1}{\gamma_{k,N}},$
where $\alpha_{k,N}$ and $\beta_{k,N}$ are defined in (3.4). Using (3.5), we see that
(4.3) $\mathcal{R}_{(\eta)}(g_{N})(z)\,=\,\sum_{k=1}^{\infty}\eta_{k}k\gamma_{k,N}z^{k},\quad z\in\mathbb{D}.$
We have
$\left|\delta_{k+1,N}-\delta_{k,N}\right|\,=\,\frac{\left|\gamma_{k+1,N}-\gamma_{k,N}\right|}{\left|\gamma_{k+1,N}\gamma_{k,N}\right|}\asymp\frac{\left|\alpha_{k+1,N}-\alpha_{k,N}\right|}{\left|\alpha_{k+1,N}\alpha_{k,N}\right|}N^{-\frac{1+\alpha}{p}}.$
This and (3.9) give
(4.4) $\left|\frac{\delta_{k+1,N}-\delta_{k,N}}{\frac{k+1}{N}-\frac{k}{N}}\right|\,\lesssim N^{1-\frac{2+\alpha}{p}},\quad N\leq k\leq 2N.$
For each $N$, let $\Psi_{N}:\mathbb{R}\rightarrow\mathbb{R}$ be defined as follows:
  • $\Psi_{N}(x)=0$, if $x\leq 0$.
  • $\Psi_{N}(x)=0$, if $x\geq 4$.
  • Between 0 and $4$, the graph of $\Psi_{N}$ is the polygonal whose vertices are the points
    $(0,0),\,\left(\frac{N}{N},\delta_{N,N}\right),\,\left(\frac{N+1}{N},\delta_{N+1,N}\right),\,\left(\frac{N+2}{N},\delta_{N+2,N}\right),\,\dots,\left(\frac{2N}{N},\delta_{2N,N}\right),\,(4,0).$
Using (4.4), we see that $\Psi_{N}$ is a Lipschitz function and that
(4.5) $\mathcal{L}(\Psi_{N})\,\lesssim N^{1-\frac{2+\alpha}{p}}.$
Using the results obtained in section 2.3 with $\Psi=\Psi_{N}$ and $n=N$ we see that
$\left|\sum_{k=0}^{4N}\Psi_{N}\left(\frac{k}{N}\right)e^{ik\theta}\right|\lesssim N^{1-\frac{2+\alpha}{p}}\frac{1}{|1-e^{i\theta}|^{2}},\quad 0<\theta<2\pi.$
Hence, if we set
$H_{N}(z)\,=\,\sum_{k=0}^{4N}\Psi_{N}\left(\frac{k}{N}\right)z^{k},\quad z\in\mathbb{D},$
we have that
(4.6) $\|H_{N}\|_{H^{p}}\,\lesssim N^{1-\frac{2+\alpha}{p}}.$
Arguing as in the proof of Theorem 1, using the Riesz projection theorem, we obtain
(4.7) $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{A^{p}_{\alpha}}\lesssim\|H_{N}\ast\mathcal{R}_{(\eta)}(g_{N})\|_{A^{p}_{\alpha}}.$
Since $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p}_{\alpha})$, we have $\|\mathcal{R}_{(\eta)}(g_{n})\|_{A^{p}_{\alpha}}\lesssim 1$ and then, using Lemma 2 and (4.6), it follows that
$\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{A^{p}_{\alpha}}\,\lesssim N^{1-\frac{2+\alpha}{p}}.$
Using Lemma 3. 1 of [29] it follows easily that $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{H^{p}}\asymp\,N^{\frac{1+\alpha}{p}}\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{A^{p}_{\alpha}}$. Then it follows that $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{H^{p}}\lesssim N^{1-\frac{1}{p}}$, which is equivalent to saying that $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
Suppose now that $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p}_{\alpha})$. Since $g_{N}\,\rightarrow\,0$, as $N\to\infty$, uniformly in compact subsets, it follows that $\mathcal{R}_{(\eta)}(g_{N})\,\rightarrow\,0$, as $N\to\infty$, in $A^{p}_{\alpha}$. This leads to $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{A^{p}_{\alpha}}\,=\operatorname{o}\left(N^{1-\frac{2+\alpha}{p}}\right)$ and to $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{H^{p}}=\,\operatorname{o}\left(N^{1-\frac{1}{p}}\right)$. Hence, $F_{(\eta)}\in\lambda^{p}_{1/p}$. $\Box$
5. Further results
We start this section obtaining a number of results about the action of the Rhaly operators on $H^{1}$.
In part (i) of Theorem 5 we proved that if $F_{(\eta)}\in\Lambda^{1}_{1}$ then $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{1})$. In our next theorem we improve this result, showing that the condition $F_{(\eta)}\in\Lambda^{1}_{1}$ actually implies that $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{1})$. We give also two conditions which are necessary for the boundedness of the Rhaly operator $\mathcal{R}_{(\eta)}$ in $H^{1}$. Before stating the theorem, let us notice that
$\Lambda^{1}_{1}\,=\,\{f\in\rm{Hol}(\mathbb{D}):f^{\prime}\in H^{1}\}.$
Also
$\beta_{1,1}(r,f)\,=\,M_{1}(r,f^{\prime}),\quad\beta_{1,1}(f)\,=\,\|f^{\prime}\|_{H^{1}}.$
Theorem 7.
Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers.
(i) If $F_{(\eta)}^{\prime}\in H^{1}$ then $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{1},\mathcal{D}^{1}_{0})$ and $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{1})$.
(ii) If   $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{1})$ then $\sum_{n=0}^{N}n|\eta_{n}|\,=\,\operatorname{O}(N)$
(iii) If  $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{1})$ then $\|\Delta_{N}(F_{(\eta)})^{\prime}\|_{H^{1}}\,=\,\operatorname{O}(\log N)$, as $N\to\infty$.
Proof. Assume that $F_{(\eta)}^{\prime}\in H^{1}$. We already know that this implies that $\mathcal{R}_{(\eta)}$ is a bounded operator from $H^{1}$ into $\mathcal{D}^{1}_{0}$.
It is well known that there are function $h$ in $H^{1}$ for which the sequence of partial sums $S_{N}h$ does not converge. However, the following is true:
If $h\in H^{1}$ then there exists a subsequence $\{S_{N_{k}}h\}$ of the sequence of partial sums of $h$ such that $S_{N_{k}}h\to f$, as $k\to\infty$, in the norm of $H^{1}$.
For instance, this can be deduced easily from the fact that $\|h_{r}-h\|_{H^{1}}\to 0$, as $r\to 1$.
So, take a subsequence $\{S_{N_{k}}(F_{(\eta)}^{\prime})\}$ of the sequence of the partial sums of $F_{(\eta)}^{\prime}$ converging to $F_{(\eta)}^{\prime}$ in $H^{1}$. Arguing as in the proof of Proposition 2 we obtain that, if $\gamma>1$ and the operators $\mathcal{R}_{N_{k+1}}$ are defined as in (3.27), then, for $f\in H^{1}$, we have
$\|\left(\mathcal{R}_{(\eta)}\,-\,\mathcal{R}_{N_{k+1}}\right)(f)\|_{\mathcal{D}^{1}_{0}}\,\lesssim\,$ $\beta_{1,1}\left(F_{(\eta)}\,-\,S_{N_{k+1}}F_{(\eta)}\right)\|f\|_{H^{1}}$
$\,=\,$ $\|F_{(\eta)}^{\prime}\,-\,\left(S_{N_{k+1}}F_{(\eta)}\right)^{\prime}\|_{H^{1}}\|f\|_{H^{1}}$
$\,=\,$ $\|F_{(\eta)}^{\prime}\,-\,S_{N_{k}}\left(F_{(\eta)}^{\prime}\right)\|_{H^{1}}\|f\|_{H^{1}}.$
Then we see that $\mathcal{R}_{(\eta)}\,-\,\mathcal{R}_{N_{k+1}}\to 0$, as $k\to\infty$ in the norm of the space $\mathcal{B}(H^{1},\mathcal{D}^{1}_{0})$. Since the operators $\mathcal{R}_{N_{k+1}}:H^{1}\,\rightarrow\,\mathcal{D}^{1}_{0}$ are finite rank operators, it follows that $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{1},\mathcal{D}^{1}_{0})$ and, then also $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{1})$. Thus (i) is proved.
Suppose now that $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{1})$. For $N\,=\,2,3,4,\dots$ set $a_{N}\,=\,1-\frac{1}{N}$ and
$f_{N}(z)\,=\,\frac{1}{N}\frac{a_{N}z}{(1-a_{N}z)^{2}}\,=\,\frac{1}{N}\sum_{n=1}^{\infty}na_{N}^{n}z^{n},\quad z\in\mathbb{D}.$
Then, for all $N$, we have that  $f_{N}\in H^{1}$ and $\|f_{N}\|_{H^{1}}\leq 1$. This implies that
(5.1) $\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{1}}\lesssim 1.$
We have
(5.2) $\mathcal{R}_{(\eta)}(f_{N})(z)=\frac{1}{N}\sum_{n=1}^{\infty}\eta_{n}\left(\sum_{k=1}^{n}ka_{N}^{k}\right)z^{n},\quad z\in\mathbb{D}.$
Using Hardy’s inequality (see section 3. 6 of [11]), we obtain
$\frac{1}{N}\sum_{n=1}^{N}\frac{|\eta_{n}|}{n+1}\left(\sum_{k=1}^{n}ka_{N}^{k}\right)\leq\frac{1}{N}\sum_{n=1}^{\infty}\frac{|\eta_{n}|}{n+1}\left(\sum_{k=1}^{n}ka_{N}^{k}\right)\leq\pi\|\mathcal{R}_{(\eta)}(f_{a})\|_{H^{1}}\,\lesssim 1.$
Since there exists a positive constant $A$ such that $a_{N}^{k}\geq A$ for $k=1,2,\dots,N$, we obtain $\frac{1}{N}\sum_{n=0}^{N}n|\eta_{n}|\lesssim 1$. Hence (ii) is proved.
Set also
$\varphi_{N}(z)\,=\,\sum_{n=N}^{2N-1}\frac{1}{\sum_{k=1}^{n}ka_{N}^{k}}z^{n}\,=\,z^{N}\psi_{N}(z),\quad z\in\mathbb{D},$
where
$\psi_{N}(z)\,=\,\sum_{n=0}^{N-1}\frac{1}{\sum_{k=1}^{n+N}ka_{N}^{k}}z^{n},\quad z\in\mathbb{D}.$
Using (5.2) we see that
$\Delta_{N}(F_{(\eta)})\,=\,N\mathcal{R}_{(\eta)}(f_{N})\ast\varphi_{N}.$
Then it follows that
(5.3) $\|\Delta_{N}(F_{(\eta)})\|_{H^{1}}\,\lesssim N\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{1}}\|\varphi_{N}\|_{H^{1}}\lesssim N\|\psi_{N}\|_{H^{1}}.$
Notice that the sequence of Taylor coefficients of $\psi_{N}$ is decreasing. Then using [33, Theorem 1. 1], we obtain
$\|\psi_{N}\|_{H^{1}}\,\asymp\,\sum_{n=0}^{N-1}\frac{1}{(n+1)\sum_{k=1}^{n+N}ka_{N}^{k}}\,\leq\,\sum_{n=0}^{N-1}\frac{1}{(n+1)\sum_{k=1}^{N}ka_{N}^{k}}\,\asymp\,\frac{\log N}{N^{2}}.$
This and (5.3) shows that $\|\Delta_{N}(F_{(\eta)})\|_{H^{1}}\,\lesssim\,\frac{\log N}{N}$. Then, using Lemma 3. 4 of [10], we obtain
$\|\Delta_{N}(F_{(\eta)})^{\prime}\|_{H^{1}}\,\lesssim\,\log N.$
Hence, (iii) is proved.
$\Box$
Let us turn to prove Theorem 4.
Proof ofTheorem 4.
If  $\eta_{n}\,=\,\operatorname{O}\left(\frac{1}{n}\right)$, as $n\to\infty$ then
$M_{2}\left(r,F_{(\eta)}^{\prime}\right)\,=\,\left(\sum_{n=1}^{\infty}n^{2}|\eta_{n}|^{2}r^{2n-2}\right)^{1/2}\,\lesssim\left(\sum_{n=1}^{\infty}r^{2n-2}\right)^{1/2}\,\lesssim\frac{1}{(1-r)^{1/2}}.$
This implies that $F_{\mu}\in\Lambda^{2}_{1/2}$ and then it follows that $F_{\mu}\in\Lambda^{q}_{1/q}$ for all $q\in[2,\infty)$. Using Theorem 1, (i) follows.
We shall use Rademacher functions and Khinchine’s inequality to prove (ii). Let $\{r_{k}\}_{k=0}^{\infty}$ be the sequence of Rademacher functions, that is,
$r_{k}(t)\,=\,\left\{\begin{array}[]{ll}1,&\hbox{if $\frac{2n}{2^{k+1}}\leq t<\frac{2n+1}{2^{k+1}}$,\,\,$n=\,0,1,2,\dots,2^{k}-1$}.\\ \\ -1,&\hbox{if $\frac{2n+1}{2^{k+1}}\leq t<\frac{2n+2}{2^{k+1}}$,\,\,\,$n=\,0,1,2,\dots,2^{k}-1$}.\end{array}\right$.
Khinchine’s inequality is the following (see, e. g., Theorem 8. 4 in chapter V of [44] or Appendix A of [11]).
Theorem B.
[Khinchine’s inequality]. For $0<p<\infty$ there exist constants $A_{p}$, $B_{p}$ with $0<A_{p}<B_{p}<\infty$ such that for all natural numbers $m$ and all $m+1$-tuples of complex numbers $c_{0},c_{1},\dots,c_{m}$, we have
$A_{p}\left(\sum_{j=0}^{m}|c_{j}|^{2}\right)^{p/2}\,\leq\int_{0}^{1}\left|\sum_{j=0}^{m}c_{j}r_{j}(t)\right|^{p}\,dt\,\leq B_{p}\left(\sum_{j=0}^{m}|c_{j}|^{2}\right)^{p/2}.$
Let $P$ is a complex polynomial, $P(z)=\sum_{j=0}^{m}a_{j}z^{j}$. We define
$P_{t}(z)\,=\,\sum_{j=0}^{m}r_{j}(t)a_{j}z^{j},\quad 0\leq t<1.$
Using Khinchine’s inequality with $c_{j}=a_{j}e^{ij\theta}$ ($0\leq j\leq m$), we obtain
$A_{p}\left(\sum_{j=0}^{m}|a_{j}|^{2}\right)^{p/2}\,\leq\int_{0}^{1}|P_{t}(e^{i\theta})|^{p}\,dt\,\leq\,B_{p}\left(\sum_{j=0}^{m}|a_{j}|^{2}\right)^{p/2},\quad\theta\in\mathbb{R}.$
This implies that
$A_{p}\left(\sum_{j=0}^{m}|a_{j}|^{2}\right)^{p/2}\,\leq\frac{1}{2\pi}\int_{0}^{2\pi}\int_{0}^{1}|P_{t}(e^{i\theta})|^{p}\,dt\,d\theta\,\leq\,B_{p}\left(\sum_{j=0}^{m}|a_{j}|^{2}\right)^{p/2}.$
Interchanging the order of integration in the second term, it follows that
$A_{p}\left(\sum_{j=0}^{m}|a_{j}|^{2}\right)^{p/2}\,\leq\int_{0}^{1}\|P_{t}\|_{H^{p}}^{p}\,dt\,\leq\,B_{p}\left(\sum_{j=0}^{m}|a_{j}|^{2}\right)^{p/2}.$
Clearly, this gives the following:
(5.4) There exists $t\in[0,1)$ such that $\|P_{t}\|_{H^{p}}\geq A_{p}\left(\sum_{j=0}^{m}|a_{j}|^{2}\right)^{1/2}.$
Suppose now that $1\leq p<2$.
Set $h(z)\,=\,\log\frac{1}{1-z}\,=\,\sum_{n=1}^{\infty}\frac{z^{n}}{n}$ ($z\in\mathbb{D}$). For all $k$, we have
$\left(\Delta_{2^{k}}h\right)^{\prime}(z)\,=\,\sum_{j=2^{k}}^{2^{k+1}-1}z^{j-1},\quad z\in\mathbb{D},\,\,\,k=0,1,2,\,\dots.$
Bearing in mind (5.4), for each $k$, let us pick $t_{k}\in[0,1)$ such that
$\|\left(\Delta_{2^{k}}h\right)^{\prime}_{t_{k}}\|\geq A_{p}2^{k/2}.$
Set
$\Upsilon(z)\,=\,\sum_{k=0}^{\infty}\left(\Delta_{2_{k}}h\right)_{t_{k}}(z),\quad z\in\mathbb{D}.$
Clearly, $\Upsilon\in\rm{Hol}(\mathbb{D})$ and $\Delta_{2^{k}}\Upsilon=\left(\Delta_{2^{k}}h\right)_{t_{k}}$, for all $k$. Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be defined by $\eta_{n}=\frac{\Upsilon^{(n}(0)}{n!}$. Then $\Upsilon=F_{(\eta)}$. We have that $|\eta_{n}|=\frac{1}{n}$ for all $n\geq 1$. Also
(5.5) $\|\left(\Delta_{2^{k}}F_{(\eta)}\right)^{\prime}\|_{H^{p}}\,=\,\|\left(\Delta_{2^{k}}h\right)^{\prime}\|_{H^{p}}\geq A_{p}2^{k/2}.$
If $1<p<2$, then (5.5) implies that $F_{(\eta)}\not\in\Lambda^{p}_{1/p}$ and then, using Theorem 1, we deduce that $\mathcal{R}_{(\eta)}\not\in\mathcal{B}(H^{p})$.
If $p=1$, then (5.5) and part (iii) of Theorem 7 show that $\mathcal{R}_{(\eta)}\not\in\mathcal{B}(H^{1})$.
Part (iii) of Theorem 4 follows immediately from the following result (see [20] for $p\geq 2$, and [30, Lemma 1] for $1<p<\infty$).
Proposition A.
If  $1<p<\infty$ and $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ is a decreasing sequence of non negative real numbers, then $F_{(\eta)}\in\Lambda^{p}_{1/p}$ if and only if $\eta_{n}=\operatorname{O}\left(\frac{1}{n}\right)$, as $n\to\infty$.
$\Box$
For $1\leq p\leq 2$ we can prove the following.
Proposition 3.
Suppose that $1\leq p\leq 2$ and let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers such that the sequence $\{|\eta_{n}|\}_{n=0}^{\infty}$ is decreasing.
If $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$ then $\eta_{n}=\operatorname{O}\left(\frac{1}{n}\right)$, as $n\to\infty$.
Proof.
If $p=1$ the result follows readily from part (ii) of Theorem 7.
Assume that $\{|\eta_{n}|\}_{n=0}^{\infty}$ is decreasing, $1<p\leq 2$, and $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$. For $N=2,3,4,\dots$, set $a_{N}=1-\frac{1}{N}$ and let $f_{N}$ be defined as in (3.1). We have
$\mathcal{R}_{(\eta)}(f_{N})(z)=\frac{1}{N^{2-\frac{1}{p}}}\sum_{n=1}^{\infty}\eta_{n}\left(\sum_{k=0}^{n}ka_{N}^{k}\right)z^{n},\quad z\in\mathbb{D},$
and $\|R_{(\eta)}(f_{N})\|_{H^{p}}\lesssim 1$. Using this, the facts that $\{|\eta_{n}|\}$ is decreasing and that there exists $A>0$ such that $a_{N}^{k}\geq A$ for $0\leq k\leq N$, and Theorem 6. 2 of [11], we deduce that
$|\eta_{N}|^{p}\frac{1}{N^{2p-1}}\sum_{n=1}^{N}(n+1)^{p-2}\left(\sum_{k=0}^{n}k\right)^{p}\lesssim\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}^{p}\lesssim 1.$
Since
$\frac{1}{N^{2p-1}}\sum_{n=1}^{N}(n+1)^{p-2}\left(\sum_{k=0}^{n}k\right)^{p}\asymp\frac{1}{N^{2p-1}}\sum_{n=1}^{N}(n+1)^{3p-2}\asymp N^{p},$
it follows that $|\eta_{N}|=\operatorname{O}\left(\frac{1}{N}\right)$. $\Box$
We remark that the sequence $\{\eta_{n}\}$ constructed to prove part (ii) of Theorem 4 provides us with an example which shows that the converse of Proposition 3 is not true for $1\leq p<2$.
Data Availability. All data generated or analyzed during this study are included in this article and in its bibliography
Conflict of interest. The authors declare that there is no conflict of interest.